OpenStudy (anonymous):

A circle has a diameter with endpoints of (-3, 8) and (7, 4). What is the center of the circle? a(10, 12) b(5, 12) c(2, 6)

OpenStudy (sriram):

it is the mid point of the 2 given points

OpenStudy (anonymous):

find the midpoint. add the x-coordinates and divide by 2. add the y-coordinates and divide by 2 C

OpenStudy (sriram):

mid point =( [-3+8]/2, [8+4]/2) =(5/2 , 12/2) =(2.5,6)

OpenStudy (anonymous):

is the queston correct?????

OpenStudy (anonymous):

idont know, is the answeer correct>?

OpenStudy (anonymous):

(-3+7)/2 = 2 (8+4)/2 = 6 (2,6) C

OpenStudy (anonymous):

SORRY!!! the ans is (2,6)